# Inferring fine-grained migration patterns across the United States

### v2

@@ 2.2 Validating MIGRATE
We show that MIGRATE estimates also increase agreement with Census datasets compared to using the Infutor data. We note that Infutor systematically undercounts the population (Table 1); to compare to Infutor as generously as possible, and in particular to ensure that MIGRATE does not reduce error due to trivial rescalings, we multiply each raw Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population prior to conducting these comparisons. Figures 2c and 2d report the reduction in error that MIGRATE estimates achieve relative to Infutor data. We compute root mean squared error (RMSE) between (1) MIGRATE estimates and ground-truth Census data and (2) Infutor data and ground-truth Census data, and report the reduction in RMSE from using MIGRATE estimates. We show average reduction in RMSE across all data releases between 2010 and 2019; error bars represent standard deviation across these releases.
MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates. Relative to raw Infutor data, MIGRATE estimates reduce error in Census tract populations by an average of $85.9\%$ and error in CBG populations by $83.6\%$. MIGRATE estimates reduce error in state-to-state flows by $89.5\%$ and error in state in-migration rates by $83.3\%$. We provide error reductions for additional metrics (movers, non-movers, and in-migration) in Appendix A.1.
@@
We again confirm that we still see these reductions in error on held-out data which is not used in our estimation procedure. Specifically, MIGRATE reduces error in county-to-county flows by $90.6\%$ and error in county in-migration rates by $49.9\%$; neither of these datasets was used in producing MIGRATE estimates. We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Appendix A.2). Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data.
@@
MIGRATE estimates reduce demographic biases present in raw Infutor data
@@
Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates. Infutor data displays biases that MIGRATE estimates greatly reduce. (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it. MIGRATE estimates remove all county-level errors by construction. (b) Spearman correlation between county demographics (x-axis) and error in Infutor estimates (y-axis). Infutor’s error is correlated with racial, socioeconomic, and other demographic characteristics. (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green). Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data. MIGRATE greatly reduces biases for all demographic subgroups.
@@
The Infutor data displays geographic and demographic biases (Figure 3). Specifically, it overrepresents populations in the counties in the northeastern United States while underrepresenting populations in the southwest (Figure 3a). [^3] For this analysis, as above, we multiply each Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population to avoid deviations due to trivial rescalings. Errors plotted are averages of county population errors across all 10 years of data from 2010-2019. MIGRATE estimates remove all such county-level errors by construction.
@@
Errors in the Infutor data also correlate with demographics (Figure 3b). Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups. These biases are consistent with biases found in previous research [22], as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
@@
MIGRATE estimates reduce biases in Infutor data. Figure 3c compares the demographic biases of the MIGRATE estimates to the biases of the Infutor data. To quantify bias, we compare (1) the ground-truth number of people within each group (from Census data) to (2) the number of people within each group estimated from Infutor or MIGRATE, which we compute as $\sum_{i}p_{i}\cdot n_{i}$, where $i$ indexes CBGs, $p_{i}$ is the proportion of people in a CBG within a given group (e.g., the proportion of Black residents) based on Census data, and $n_{i}$ is the number of people in a CBG in Infutor or MIGRATE. This quantifies how much Infutor and MIGRATE overcount or undercount a given demographic group, relative to ground-truth Census data. The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases. We present analyses for additional demographic groups, finding similar results, in Appendix A.4.
2.3 Analysis of national migration patterns
@@
Figure 4: National migration statistics. (a) Flows between ten types of CBGs — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile. Rows correspond to the origin CBG, and columns to the destination CBG; for example, the top left entry indicates that 90% of movers from plurality white CBGs move to plurality white CBGs. The final two rows report the proportion of all movers moving to CBGs of each type, and the population share living in CBGs of each type. We report averages across all years. (b) Probability of moving to a higher-median-income CBG, conditional on income decile of origin CBG, and plurality race of origin CBG. (c) Distance moved stratified by CBG type.
@@
We use MIGRATE to analyze national patterns of migration from 2010-2019. The spatial granularity of the data affords a unique opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation. We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories [^4] To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates. For example, to classify CBGs into plurality races to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates..
Figure 4a plots flows between the ten categories — for example, the proportion of movers from top-income-quartile CBGs who move to top-income-quartile CBGs. This reveals substantial homophily in migration: out-movers from all ten CBG categories are more likely than movers as a whole to move to CBGs of the same category. However, the strength of this homophily varies across categories. Movers from plurality Black, Asian, and Hispanic CBGs exhibit strong homophily: they are 5.5$\times$, 14.6$\times$, and 4.3$\times$ likelier than movers as a whole to move to CBGs with the same plurality race group (Figure S3a reports relative rates for all groups). We also observe income homophily, particularly for movers within top and bottom quartile CBGs: movers from bottom-income-quartile CBGs are nearly twice as likely as movers as a whole to move to CBGs in the bottom income quartile (34% versus 18% for movers as a whole); movers from top-income-quartile CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs (53% versus 32%). In Figure S3b, we confirm that this homophily does not occur merely because many moves are local and demographics are spatially correlated: we also observe homophily when restricting to long-distance (out-of-county) moves.
Figure 4a also shows that people are likelier to move to CBGs in higher income quartiles: 32% of moves are to top-income-quartile CBGs, while only 18% are to bottom-income-quartile CBGs. This trend becomes more pronounced over the decade we study (Figure S4d) and is only partially explained by the fact that top-income-quartile CBGs account for a larger share of the population (29%) than bottom-income-quartile CBGs (21%). There are also racial disparities: movers from plurality Asian CBGs are 1.7$\times$ likelier than movers as a whole to move to top-income-quartile CBGs; movers from plurality Black CBGs are 2.0$\times$ likelier than movers as a whole to move to bottom-income-quartile CBGs.
@@
Figure 4b investigates whether these racial disparities persist when controlling for origin CBG median income: in particular, plotting the share of movers moving to a higher-income CBG when controlling for the mover’s origin CBG median income decile. This reveals that the probability of moving to a higher-income CBG varies substantially by race even conditional on origin CBG income: movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs. For example, movers from fifth-income-decile, plurality Asian CBGs have a 71% chance of moving to a higher-income CBG; movers from fifth-income-decile, plurality Black CBGs have a 53% chance. Figure S4 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG. Overall, we find robust and substantial racial disparities in income of destination CBG that persist even conditional on income of origin CBG.
@@
Finally, Figure 4c provides statistics on the distance of moves. 37% of movers move less than 5 miles; 40% from 5-50 miles; and 23% more than 50 miles. Figure 4c also highlights demographic variation in these statistics, revealing that movers from plurality white CBGs, rural CBGs, and higher-income CBGs are likelier to move long distances (more than 50 miles). Appendix B.4 reports additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and shows that migration distance increases over the decade we study.
Overall, these results demonstrate that fine-grained migration data illuminates important demographic variation in homophily, upward mobility, and migration distance.
2.4 Analysis of local migration patterns
Figure 5: Migration in response to California wildfires. (a) Out-migration following the Camp fire (2018; top) and Tubbs fire (2017; bottom). Red boundaries plot fire perimeters; black lines plot county boundaries; CBGs are colored by domestic out-migration rate in MIGRATE. Out-migration rates exceed 50% in many CBGs within the fire perimeters. (b) Out-migration rates in different groups of CBGs over time according to MIGRATE estimates. Out-migration rates in the year after the fire are higher in CBGs within the fire perimeter (red line) than in groups of CBGs outside the fire perimeter (other lines), including those neighboring the fire perimeter, those in affected counties, or those within California. (c) Out-migration rates in ACS 5-year county-to-county data remain relatively constant over time.
@@ 2.4 Analysis of local migration patterns
In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California. Natural disasters, including wildfires, are known drivers of human mobility [3, 29]. There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area. In many US states, including California, wildfire risk increased from 2010-2019 [30], and is expected to further increase due to climate change [31]. Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters [32].
@@
Post-wildfire migration estimates can inform disaster response and long-term planning in multiple ways. First, migration data is crucial for policymaking in the aftermath of a wildfire. It helps guide the allocation of housing, public health, and other resources to support displaced residents while minimizing strain on housing markets in receiving areas [33, 29, 34, 35]. Data on out-migration rates from affected areas can also inform decisions about whether and how to rebuild in fire-prone regions [36, 37]. Second, high-resolution migration data is important for future wildfire planning. As populations relocate in response to shifting wildfire risks, migration patterns can refine policymakers’ estimates of future risk and guide the regulation of housing and insurance markets [38, 39, 40, 41, 42]. While some households may move to reduce their exposure, others may become “trapped” in high-risk areas due to financial, social, or logistical barriers [43, 44, 45]. Identifying these communities through high-resolution migration data can help ensure that government support reaches those who need it most [46, 37].
@@
We use high-resolution fire perimeter data from the California Department of Forestry and Fire Protection [47]. Fire perimeters are typically much smaller than county boundaries, suggesting that analyzing fire impacts may require the granularity of the MIGRATE estimates as opposed to the relatively coarse county-to-county flows. We analyze the two most destructive fires in California from 2010-2019 (as quantified by the number of structures lost to the fire): the Tubbs Fire in October 2017, and the Camp Fire in November 2018 (Figure 5). The Tubbs Fire occurred in Napa County and Sonoma County and destroyed at least 5,636 residential or commercial structures [48]. The Camp Fire, about a year later, destroyed over 18,804 structures and damaged nearly 1,000 more in the Northern California county of Butte [49] [^5] The Camp Fire remains the most destructive wildfire in California as of early 2025; the Tubbs Fire has only been surpassed by the January 2025 Eaton and Palisades fires [50]..
MIGRATE estimates reveal dramatic levels of out-migration for CBGs which lie within the fire perimeters (Figure 5a) in the year following the fires, often exceeding 50%. [^6] The fact that the out-migration rate is not 100% is likely due to a number of factors, including the often-preferred option of rebuilding and returning [40, 41]. Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [32]. In contrast, CBGs that do not lie within the fire perimeters experience much lower out-migration rates. We systematically quantify these differences in Figure 5b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole. The out-migration rate in the year following the Camp Fire for CBGs within the fire perimeter is 46%, at least 3.1$\times$ that of less-affected CBG groups; the out-migration rate in the year following the Tubbs Fire for CBGs within the perimeter is 37%, at least 2.8$\times$ that of less-affected CBG groups. Our estimates of out-migration following the Camp Fire are similar to those in prior work [29].
@@
In Appendices C and D, we provide two additional analyses of local migration patterns that are impossible using only county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy. Collectively, these analyses showcase how MIGRATE estimates reveal important, policy-relevant local migration patterns that traditional data sources conceal.

### v3

@@ 3.2 MIGRATE estimates reduce error relative to raw Infutor data
We show that MIGRATE estimates also increase agreement with Census datasets compared to using the Infutor data alone. We note that Infutor systematically undercounts the population, as shown in Table 1; to compare to Infutor as generously as possible, and in particular to ensure that MIGRATE does not reduce error due to trivial rescalings, we multiply each raw Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population prior to conducting these comparisons. Figures 2d-f report the reduction in error that MIGRATE estimates achieve relative to Infutor data. We compute root mean squared error (RMSE) between (1) MIGRATE estimates and ground-truth Census data and (2) Infutor data and ground-truth Census data, and report the reduction in RMSE from using MIGRATE estimates. We report average reduction in RMSE across all data releases between 2010 and 2019; error bars represent standard deviation across these releases.
@@
MIGRATE estimates eliminate error in state and county populations because they are constructed to match county-level estimates. They also reduce error in Census Tract and CBG populations by an average of $85.9\%$ and $83.6\%$ respectively (Figure 2d); in state-to-state movers and county-to-county movers by $87.5\%$ and $42.3\%$ (Figure 2e); and in state in-migration and county in-migration rates by $87.3\%$ and $51.8\%$ (Figure 2f). These gains persist even when using held-out datasets that are not used in estimating MIGRATE (namely, county-to-county flows and county in-migration rates). We also repeat the validation above where we successively remove types of data (e.g., CBG-level populations) from our estimation procedure, and verify that our estimation procedure still reduces error in reproducing those held-out data types (Supplementary Table 1(b)).
Collectively, these validations demonstrate that MIGRATE estimates increase agreement with Census datasets, including held-out datasets, compared to raw Infutor data. We provide error reductions for additional metrics (all flows, non-movers, and in-migration counts as opposed to rates) in Supplementary Table 1(b), and conduct additional validations on synthetic data (Appendix A.2).
@@
3.3 MIGRATE estimates reduce demographic biases present in raw Infutor data
@@
Figure 3: Assessment of demographic bias in the raw Infutor data and the MIGRATE estimates. Infutor data displays biases that MIGRATE estimates greatly reduce. (a) Average errors in county populations in Infutor data relative to Census data; orange denotes counties where Infutor underrepresents the population, and purple denotes counties where Infutor overrepresents it. MIGRATE estimates remove all county-level errors by construction. (b) Spearman correlation between county demographics (x-axis) and error in Infutor estimates (y-axis). Infutor’s error is correlated with racial, socioeconomic, and other demographic characteristics. (c) Comparison of demographic bias in Infutor (purple) and MIGRATE (green). Bars compare demographic subpopulations estimated from Infutor or MIGRATE to ground-truth Census data, averaged over $n=5$ population data releases (American Community Survey 5-year estimates, 2015 through 2018) Error bars represent standard deviations across these releases. MIGRATE greatly reduces biases for all demographic subgroups.
@@
The Infutor data displays geographic and demographic biases (Figure 3). Specifically, it overrepresents populations in the counties in the northeastern United States while underrepresenting populations in the southwest (Figure 3a) [^5] For this analysis, as above, we multiply each Infutor matrix $E^{(t)}$ by a scaling factor so it matches the national yearly Census population to avoid deviations due to trivial rescalings.. Errors plotted are averages of county population errors across all 10 years of data from 2010-2019. MIGRATE estimates remove all such county-level errors by construction.
@@
Errors in the Infutor data also correlate with demographics (Figure 3b). Infutor data overrepresents White, home-owning, richer, and older populations: there is a positive Spearman correlation between the relative error in county population estimates and the population share of a county in each of these groups. These biases are consistent with biases found in previous research ramiller_residential_2024, as we discuss in detail in Appendix A.3, and could propagate into biases in downstream analyses of inequality and other topics.
@@
MIGRATE estimates reduce biases in Infutor data. Figure 3c compares the demographic biases of the MIGRATE estimates to the biases of the Infutor data. To quantify bias, we compare (1) the ground-truth number of people within each group (from Census data) to (2) the number of people within each group estimated from Infutor or MIGRATE, which we compute as $\sum_{i}p_{i}\cdot n_{i}$, where $i$ indexes CBGs, $p_{i}$ is the proportion of people in a CBG within a given group (e.g., the proportion of Black residents) based on Census data, and $n_{i}$ is the number of people in a CBG in Infutor or MIGRATE. This quantifies how much Infutor and MIGRATE overcount or undercount a given demographic group, relative to ground-truth Census data. The raw Infutor data substantially overcounts rural, older, white, and home-owning populations, and undercounts younger, Black, Asian, Hispanic, renter, and below-the-poverty-line populations; the MIGRATE estimates almost entirely eliminate these biases. We present analyses for additional demographic groups, including immigrant populations, in Appendix A.4.
4 Analysis of national migration patterns
@@
Figure 4: National migration statistics. MIGRATE can be used to reveal fine-grained patterns of mobility across the United States. (a) Flows between ten types of Census Block Groups (CBGs) — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile. Rows correspond to the origin CBG, and columns to the destination CBG; for example, the top left entry indicates that 90% of movers from plurality white CBGs move to plurality white CBGs. The final two rows report the proportion of all movers moving to CBGs of each type, and the population share living in CBGs of each type. We report averages across all years. (b) Probability of moving to a higher-median-income CBG, conditional on income decile of origin CBG, and plurality race of origin CBG. (c) Distance moved stratified by CBG type.
@@
We use MIGRATE to analyze national patterns of migration from 2010-2019. The spatial granularity of the data affords an opportunity to study demographic variation in migration patterns — for example, differences in migration between higher-income and lower-income CBGs — that may be obscured by county or state-level data, which show far less demographic variation. We hence divide CBGs into 10 (overlapping) categories — plurality white, Asian, Black, and Hispanic; urban versus rural; and bottom, second, third, and top income quartile — and stratify the migration statistics we compute by these ten categories. To determine the category of each CBG every year, we use the most recent ACS 5-year demographic estimates. For example, to classify CBGs by their plurality race group to study moves in the 2010-2011 period, we use the ACS 2006-2010 race and ethnicity estimates.
[…]
@@
Figure 4b investigates whether these racial disparities persist when controlling for origin CBG median income: in particular, plotting the share of movers moving to a higher-income CBG when controlling for the mover’s origin CBG median income decile. This reveals that the probability of moving to a higher-income CBG varies substantially by race even conditional on origin CBG income: movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs. For example, movers from fifth-income-decile, plurality Asian CBGs have a 71% chance of moving to a higher-income CBG; movers from fifth-income-decile, plurality Black CBGs have a 53% chance. Supplementary Figure 5 shows that the same racial disparities occur when controlling for CBG median income percentile, as opposed to decile; or when examining the probability of moving to a top- or bottom-income-quartile CBG, as opposed to moving to a higher-income CBG. Overall, we find robust and substantial racial disparities in income of destination CBG that persist even conditional on income of origin CBG.
@@
Finally, Figure 4c provides statistics on the distance of moves. 37% of movers move less than 5 miles; 40% from 5-50 miles; and 23% more than 50 miles. Figure 4c also highlights demographic variation in these statistics, revealing that movers from plurality white CBGs, rural CBGs, and higher-income CBGs are likelier to move long distances (more than 50 miles). In Appendix B.4, we report additional migration distance statistics stratified by geographic boundary (i.e., whether movers remain within the same tract, county, or state) and show that migration distance increases over the decade we study.
Overall, these results demonstrate that fine-grained migration data illuminates important demographic variation in homophily, upward mobility, and migration distance. Future research could stratify migration flows by additional characteristics available in Census data: for example, one might study the migration patterns of immigrants by analyzing flows from areas with larger proportions of immigrants, or larger proportions of residents with a given country of birth—both of these are datasets available at the sub-county level via ACS. However, we note that such analyses require significant heterogeneity across CBGs (or other Census areas) in the demographic trait being studied; for example, it would not be possible to reliably disaggregate migration patterns by gender using this method, since gender proportions remain relatively stable across Census areas.
5 Analysis of local migration patterns
@@
In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California. Natural disasters, including wildfires, are known drivers of human mobility mcconnell_rare_2024, mcconnell2021effects. There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area. In many US states, including California, wildfire risk increased from 2010-2019 modaresi_rad_human_2023, and is expected to further increase due to climate change Barbero_2015. Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters din_administrative-data.
Figure 5: Migration in response to California wildfires. (a) Out-migration following the Camp fire (2018; top) and Tubbs fire (2017; bottom). Red boundaries plot fire perimeters; black lines plot county boundaries; Census Block Groups (CBGs) are colored by domestic out-migration rate in MIGRATE. Out-migration rates exceed 50% in many CBGs within the fire perimeters. (b) Out-migration rates in different groups of CBGs over time according to MIGRATE estimates. Out-migration rates in the year after the fire are higher in CBGs within the fire perimeter (red line) than in groups of CBGs outside the fire perimeter (other lines), including those neighboring the fire perimeter, those in affected counties, or those within California. (c) Out-migration rates in the American Community Survey (ACS) 5-year county-to-county data remain relatively constant over time.
@@
Post-wildfire migration estimates can inform disaster response and long-term planning in multiple ways. First, migration data is crucial for policymaking in the aftermath of a wildfire. It helps guide the allocation of housing, public health, and other resources to support displaced residents while minimizing strain on housing markets in receiving areas rosenthal2021health, mcconnell2021effects, hopfer2024repeat, isaac2023differences. Data on out-migration rates from affected areas can also inform decisions about whether and how to rebuild in fire-prone regions mcwethy2019rethinking, mach2021reframing. Second, high-resolution migration data is important for future wildfire planning. As populations relocate in response to shifting wildfire risks, migration patterns can refine policymakers’ estimates of future risk and guide the regulation of housing and insurance markets boomhower_how_2024, baylis_adapatation_2024, kramer_post-wildfire_2021, alexandre_rebuilding_2014, lee_post-disaster_2024. While some households may move to reduce their exposure, others may become “trapped” in high-risk areas due to financial, social, or logistical barriers bohra2014nonlinear, cattaneo2016migration, benveniste2024global. Identifying these communities through high-resolution migration data can help ensure that government support reaches those who need it most hino2017managed, mach2021reframing.
@@
We use high-resolution fire perimeter data from the California Department of Forestry and Fire Protection california_department_of_forestry_and_fire_protection_historical_2024. Fire perimeters are typically much smaller than county boundaries, suggesting that analyzing fire impacts may require the granularity of the MIGRATE estimates as opposed to the relatively coarse county-to-county flows. We analyze the two most destructive fires in California from 2010-2019 (as quantified by the number of structures lost to the fire): the Tubbs Fire in October 2017, and the Camp Fire in November 2018 (Figure 5). The Tubbs Fire occurred in Napa County and Sonoma County and destroyed at least 5,636 residential or commercial structures california_department_of_forestry_and_fire_protection_tubbs_2025. The Camp Fire, about a year later, destroyed over 18,804 structures and damaged nearly 1,000 more in the Northern California county of Butte california_department_of_forestry_and_fire_protection_camp_2025. The Camp Fire remains the most destructive wildfire in California as of early 2025; the Tubbs Fire has only been surpassed by the January 2025 Eaton and Palisades fires california_department_of_forestry_and_fire_protection_top_2025.
[…]
@@
In Appendix C, we provide two additional analyses of local migration patterns that are impossible using county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy. Collectively, these analyses showcase how MIGRATE estimates reveal important, policy-relevant local migration patterns that traditional data sources conceal.